### v2

# Metric Oja depth, new statistical tool for estimating the most central objects

Abstract
The Oja depth (simplicial volume depth) is one of the classical statistical techniques for measuring the central tendency of data in multivariate space. Despite the widespread emergence of object data like images, texts, matrices or graphs, a well-developed and suitable version of Oja depth for object data is lacking. To address this shortcoming, a novel measure of statistical depth, the metric Oja depth applicable to any object data, is proposed. Two competing strategies are used for optimizing metric depth functions, i.e., finding the deepest objects with respect to them. The performance of the metric Oja depth is compared with three other depth functions (half-space, lens, and spatial) in diverse data scenarios.
keywords:
Object data , Metric Oja depth , Statistical depth , Optimization , Metric statistics
[…]
[label1]organization=Department of Mathematics and Statistics,
University of Turku,country=Finland
1 Introduction
1.1 Background
Nowadays, data is being generated in high volumes, at high speeds, and with great diversity. A key aspect of modern data analysis is that we no longer deal exclusively with data existing in Euclidean spaces. Instead, data are currently taking on more complex formats, such as images, graphs, matrices, etc., all of which reside in non-Euclidean spaces and are collectively known as object data. Methodology for object data, known as object data analysis or metric statistics [8], are currently bountiful in statistical literature, see, e.g., [9, 29, 30, 31]. Object data are typically modelled as residing in an arbitrary metric space $(\mathcal{X},d)$ and we also adopt this approach.
As with any data, the analysis of object data begins with exploratory data analysis on the raw data. This can take various forms: calculating descriptive statistics, dimension reduction, detecting outliers etc. However, despite being straightforward for Euclidean data, these tasks can be surprisingly difficult in the context of object data, due to the lack of structure in $(\mathcal{X},d)$. In this work, we focus on one of the most fundamental exploratory statistical tasks, location/mean/average estimation, using depth functions as our tool of choice.
The foundation for depth functions was introduced in 1975 by Tukey who proposed his seminal half-space depth for multivariate data to rank observations and reveal features of the underlying data distribution [25]. Apart from this, numerous depth functions for data in Euclidean space have been proposed. E.g., the convex hull peeling depth (onion depth) [1, 12], simplicial volume depth (Oja depth) [22] and several others, see Table A.1 in Appendix A and the comprehensive addressing in [19].
In recent years, object/metric versions of various depth functions $D$ have been proposed, see Table A.1. That is, given an object $X\in\mathcal{X}$ and a distribution $P$ taking values in $\mathcal{X}$, the depth $D(X;P)$ describes how central the object $X$ is w.r.t. $P$, much in the same way as for the Euclidean depths earlier. Two key properties of these extensions are that (a) they depend on the data only through the metric $d$, making them applicable to any form of object data, regardless in which specific metric space they live, and (b) if $(\mathcal{X},d)$ is a Euclidean space, then the original Euclidean version of the depth is recovered, showing that these metric depth functions are “true” generalizations of their classical counterparts. Metric lens depth was developed in [16, 5, 14], metric half-space depth was proposed in [7] and metric spatial depth in [27].
Most of the depth functions in Table A.1, in particular all three that have object versions, are robust [18], meaning that their performance is not skewed by the presence of possible outliers. In Section 2 we list the robustness properties of the existing metric depth functions, and more details on the robustness of Euclidean depths can be found in [21, Table 1] and [32, 20]. Note that not all location estimation procedures are robust. For example, the Fréchet mean [13] which is a generalization of the concept of average and a classical estimator of location for object data, is not robust, but is instead affected greatly by data outliers. Despite the popularity of Fréchet mean and similar tools, in this work we concentrate solely on robust methods. This is because object data is very diverse and recognizing any possible outliers in them can, as such, be very difficult. To ensure that our analyses stay reliable, we therefore find it imperative to develop robust methods, which automatically protect us against outliers, regardless of their exact type and whether we recognize them or not.
1.2 Our Contributions
This work focuses on the robust location estimation of object data using depth functions, and its main contributions are as follows.
(i) We propose a new depth function for object data, the metric Oja depth, and extensively study its theoretical properties, in particular characterizing its range of values and showing the sample consistency of its maximizers under specific assumptions.
(ii) We study the estimation of the deepest objects (data location). That is, given a sample $X_{1},\ldots,X_{n}$ of object data, we aim to find the object $X\in\mathcal{X}$ which has the maximal depth value w.r.t. the sample. Despite the fundamental nature of this problem, in the earlier works on metric depths it has been considered only by [7], who proposed finding the optimum among the sample points $X_{1},\ldots,X_{n}$, or using more elaborate manifold optimization to find the deepest out-of-sample object, as implemented in the R-package MHD [6]. As a competing approach, we propose using non-linear Euclidean optimizers along with a PCA-reduced coordinate representation of the object data to estimate the deepest out-of-sample object.
(iii) We conduct extensive simulation and data comparisons between our proposed depth and its competitors in estimating the deepest object in several data scenarios. In particular, we investigate how much the in-sample estimation can be improved by using non-linear optimization to achieve out-of-sample estimation. The results show that, at the cost of greater computational complexity, metric Oja depth surpasses its competitors in several scenarios, especially when combined with our proposed optimization strategies.
1.3 Contents
We review the three existing metric depth functions in Section 2. In Section 3, we introduce the proposed new metric depth function and develop its theoretical properties. In Section 4, we evaluate and compare it with three other depth functions, using two simulation scenarios. In Section 5, we further test our metric depth function on a real dataset, relying on computational techniques to conduct statistical inference. Finally, future recommendations and conclusions are compiled in Section 6.
2 Review of existing metric depth functions
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and take $(\mathcal{X},d)$ to be a complete and separable metric space where our data reside. Recall that a metric $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ quantifies distances between objects residing in the set $\mathcal{X}$. Formally, metrics are required to satisfy the following axioms: Positivity, $d(x,y)\geq 0$ with equality if and only if $x=y$; Symmetry, $d(x,y)=d(y,x)$; and Triangle inequality, $d(x,y)\leq d(x,z)+d(z,y)$ [3, 15]. Further, we let $P$ be a probability measure defined on the Borel sets of $\mathcal{X}$. Given a fixed, non-random object $x\in\mathcal{X}$, the metric depth functions answer the question “how central is the object $x$ with respect to the distribution PP?”.
To tackle this question, [16, 5, 14] proposed metric lens depth, defined as
$D_{L}(x):=\mathbb{P}\left(d(X_{1},X_{2})>\max\left\{d(X_{1},x),d(X_{2},x)\right\}\right),$
where $X_{1},X_{2}$ are independent random variables with the distribution $P$. Drawing an analogy from Euclidean statistics, $D_{L}(x)$ essentially gives the probability that the side corresponding to $X_{1},X_{2}$ is the longest in a “triangle” drawn using the objects $x,X_{1},X_{2}$. Intuitively, the probability is small (large) if $x$ is located far away from (close to) the bulk of the distribution $P$, making $D_{L}$ behave as expected from a depth function.
[7] gave a similar treatment to the classical half-space depth and defined the metric half-space depth as
$D_{H}(x)$ $=\inf_{{x_{1},x_{2}}\in\mathcal{X},\ d(x_{1},x)\leq d(x_{2},x)}\mathbb{P}(d(X,x_{1})\leq d(X,x_{2})).$
Every pair of objects $x_{1},x_{2}$ in the infimum divides the space $\mathcal{H}$ in two subsets (objects closer to $x_{1}$ than $x_{2}$ and vice versa). The depth $D_{H}(x)$ is defined to be the smallest possible probability mass of a subset produced in this way and containing the object $x$. As with the metric lens depth, it is easy to see that, for an outlying object $x$, it is easy to find a subset which contains $x$ but has very little $P$-probability mass, making $D_{H}(x)$ small. An approximative algorithm, that we also use in this paper to compute the metric half-space depth, is given in [7, Algorithm 1].
The metric spatial depth was proposed in [27] as
where $X_{1},X_{2}\sim P$ are independent. The interpretation of $D_{S}(x)$ is trickier than for the previous two depths, but it essentially measures how likely $x$ and two random objects $X_{1},X_{2}$ are to yield equality in the triangle inequality. The depth takes values in $[0,2]$ and the value 0 is reached if and only if the probability of equality is 1 and $x$ is never in the middle of the other two objects, whereas to reach the value $2$, $x$ must always reside between the two other objects. More details on interpreting the metric spatial are given in [27].
Throughout this work, we refer to metric lens depth, metric half-space depth and metric spatial depth as MLD, MHD and MSD, respectively. All of these have the property of vanishing at infinity meaning that, for objects far enough away from the bulk of the distribution $P$, the depth values approach zero. All three also satisfy specific forms of continuity (small changes in $x$ entail small changes in the depth) and both MLD and MHD are invariant to data transformations that preserve the ordering of distances. Regarding robustness properties, MHD and MSD are robust, the former in the sense of having a high breakdown point and the latter in the sense of having a bounded influence function. Whereas, MLD is also likely to be highly robust, being based only on comparisons between distances and not their absolute sizes, but this aspect of MLD has not yet been studied in the literature. For detailed statements of these, and some other properties of the three depth functions, including sample consistency results, we refer the reader to the papers [5, 14, 7, 27].
3 New depth for object data
3.1 Population-level properties
For any three objects $x_{1},x_{2},x_{3}\in\mathcal{X}$, we use the notation $L(x_{1},x_{2},x_{3})$ to denote the event that $d(x_{1},x_{3})=d(x_{1},x_{2})+d(x_{2},x_{3})$. On an intuitive level, $L(x_{1},x_{2},x_{3})$ means that $x_{2}$ resides “between” $x_{1}$ and $x_{3}$, since under this event the distance from $x_{1}$ to $x_{3}$ is the same as when taking a detour through $x_{2}$. Hence, in the sequel we use phrases such as $x_{2}$ lies in between $x_{1}$ and $x_{3}$” to indicate that $L(x_{1},x_{2},x_{3})$ holds. For arbitrary objects $x_{0},x_{1},x_{2},x_{3}\in\mathcal{X}$, we denote by $B_{3}(x_{0},x_{1},x_{2},x_{3})$ the $3\times 3$ matrix whose $(k,\ell)$-element equals
$\frac{1}{2}\left\{d^{2}(x_{0},x_{k})+d^{2}(x_{0},x_{\ell})-d^{2}(x_{k},x_{\ell})\right\}.$ (1)
Letting $X_{1},X_{2},X_{3}\sim P$ be i.i.d. random objects, we propose to measure the depth of a fixed object $x\in\mathcal{X}$ with respect to the distribution $P$ as
$D_{O3}(x):=\frac{1}{1+\mathrm{E}\left[\left\{|B_{3}(x,X_{1},X_{2},X_{3})|+4d^{2}(x,X_{1})d^{2}(x,X_{2})d^{2}(x,X_{3})\right\}^{1/2}\right]}.$ (2)
The subscript O refers to “Oja” due to the close connection between $D_{O3}(x)$ and the classical Oja depth explained in Appendix B. By Theorem C.1 in Appendix C, taking the square root in the denominator of (2) is well-defined. The intuitive meaning of $D_{O3}(x)$ is not clear from its definition (2) alone, so we next present a collection of results that allow better understanding it, starting with a moment condition that guarantees its existence.
Theorem 1.
Assume that, for some $a\in\mathcal{X}$, we have $\mathrm{E}\{d(X,a)\}<\infty$. Then $D_{O3}(x)$ exists as well-defined.
By “well-defined” in Theorem 1 we mean that the expected value used to compute $D_{O3}(x)$ exists as finite. The condition $\mathrm{E}\{d(X,a)\}<\infty$ is analogous to requiring a univariate random variable to have a finite mean and is a quite mild condition as even the Frechét mean already requires the existence of second moments. The choice of $a$ in Theorem 1 is completely arbitrary as, if the condition $\mathrm{E}\{d(X,a)\}<\infty$ holds for some $a$, then, by the triangle inequality, it holds for all $a\in\mathcal{X}$.
Theorems 2 and 3 next establish the range of $D_{O3}(x)$ to be $[0,1]$ and give interpretations to these endpoints.
Theorem 2.
$D_{O3}(x)$ takes values in $[0,1]$. Moreover, if $D_{O3}(x)=1$, then
$P\{L(X_{1},x,X_{2})\cup L(X_{2},x,X_{3})\cup L(X_{3},x,X_{1})\}=1,$
where $X_{1},X_{2},X_{3}\sim P$ are i.i.d.
Theorem 2 implies that large values of $D_{03}(x)$ are indicative of the centrality of the object $x$ with respect to the distribution $P$. That is, under the extremal case $D_{03}(x)=1$ the object $x$ must almost surely lie in between of at least two of three randomly sampled objects $X_{1},X_{2},X_{3}$ from $P$.
We say that a sequence of objects $x_{n}\in\mathcal{X}$ is divergent if there exists $a\in\mathcal{X}$ such that $d(x_{n},a)\rightarrow\infty$ when $n\rightarrow\infty$. Intuitively, a divergent sequence is such that it moves “towards infinity” eventually getting arbitrarily far from any fixed object. Clearly, such sequences do not exist in every metric space.
Theorem 3.
Assume that, for some $a\in\mathcal{X}$, we have $\mathrm{E}\{d^{2}(X,a)\}<\infty$. Let $x_{n}$ be a divergent sequence of points in $\mathcal{X}$. Then, $D_{O3}(x_{n})\rightarrow 0$ as $n\rightarrow\infty$.
The second moment condition in Theorem 3 is used to avoid certain pathological behavior, see the proof of the result in Appendix C. Theorem 3 states that small values of $D_{O3}(x)$ are reached by outlying objects, i.e., points that are located far away from the bulk of the distribution $P$.
It is typical for depth functions to satisfy various invariance properties [7, 14]. Namely, one usually expects that the depth of a point $x$ w.r.t. to distribution $P$ does not change if $x$ and $P$ are both subjected to a specific class of transformations. Being based on pair-wise distances, $D_{O3}$ inherits such properties directly from the metric $d$. That is, if $d(gx_{1},gx_{2})=d(x_{1},x_{2})$ for all $g\in\mathcal{G}$ where $\mathcal{G}$ is a group of transformations on $\mathcal{X}$, then also $D_{O3}$ is invariant to $\mathcal{G}$. In particular, if $x_{0}\in\mathcal{X}$ has maximal depth w.r.t. $P$, then $gx_{0}$ has maximal depth w.r.t. to $P_{g}$, the distribution of $gX$ where $X\sim P$.
3.2 Sample-level properties
Let $X_{1},\ldots,X_{n}$ be a random sample of objects from $P$. To estimate depths in practice, a natural sample counterpart of $D_{O3}$ is the third-order $U$-statistic,
$D_{O3,n}(x):=\frac{1}{1+\frac{1}{\binom{n}{3}}\sum_{(i,j,k)}h(x,X_{i},X_{j},X_{k})},$
where the summation runs over all unordered triples $(i,j,k)$ of indices from $\{1,\ldots,n\}$ with distinct elements and
$h(x,X_{1},X_{2},X_{3}):=\{|B_{3}(x,X_{1},X_{2},X_{3})|+4d^{2}(x,X_{1})d^{2}(x,X_{2})d^{2}(x,X_{3})\}^{1/2}.$
Standard results on $U$-statistics [17] with the continuous mapping theorem and Theorem 1 gives the following point-wise convergence result which states that the depth of any single object $x$ can be accurately estimated using $D_{O3,n}$.
Theorem 4.
Assume that, for some $a\in\mathcal{X}$, we have $\mathrm{E}\{d(X,a)\}<\infty$. Then, for any fixed $x\in\mathcal{X}$, we have $D_{O3,n}(x)\rightarrow_{p}D_{O3}(x)$ as $n\rightarrow\infty$.
A stronger result is obtained if the metric space is totally bounded, in which case the convergence is uniform in $\mathcal{X}$.
Theorem 5.
Assume that the metric space $(\mathcal{X},d)$ is totally bounded. Then,
$\sup_{x\in\mathcal{X}}|D_{O3,n}(x)-D_{O3}(x)|\rightarrow_{p}0,$
as $n\rightarrow\infty$.
The classical $M$-estimator argument [26, Theorem 5.7] then shows that the point with maximal depth, assuming that it is unique, can be consistently estimated with the sample depth function $D_{O3,n}$.
Corollary 6.
Assume that the metric space $(\mathcal{X},d)$ is totally bounded and that $x\mapsto D_{O3}(x)$ has unique maximizer $x_{0}$. Then any sequence $x_{n}\in\mathcal{X}$ of maximizers of $x\mapsto D_{O3,n}(x)$ satisfies $x_{n}\rightarrow_{p}x_{0}$ as $n\rightarrow\infty$.
The sample maximizers in Corollary 6 exist since $D_{O3,n}$ is a continuous function over a totally bounded and complete, and hence compact, set. The assumption of a unique population maximizer is discussed further in Section 6. Also MLD and MHD satisfy analogous consistency properties, see [5, 14, 7].
3.3 Auxiliary depth-like quantity
Our proposed depth $D_{O3}$ is based on the 3-dimensional Oja depth, see Appendix B for details, and a natural question is whether also the $p$-dimensional Oja depth can be used for similar purposes with $p\neq 3$. In this section we briefly investigate the version with $p=2$. Letting $B_{2}(x_{0},x_{1},x_{2})$ denote the $2\times 2$ top left sub-matrix of $B_{3}(x_{0},x_{1},x_{2},x_{3})$, we define
$D_{O2}(x):=\frac{1}{1+\mathrm{E}\left\{|B_{2}(x,X_{1},X_{2})|^{1/2}\right\}},$ (3)
which, as shown in Appendix B, is a metric generalization of the classical 2-dimensional Oja depth. As with $D_{O3}$, also $D_{O2}$ is well-defined as soon as the first moment exists, i.e., $\mathrm{E}\{d(X,a)\}<\infty$ for some $a\in\mathcal{X}$ (proof omitted). However, the next result, which follows directly from Theorem C.1(i) in Appendix C shows that $D_{O2}(x)$ does not actually measure the centrality of $x$ w.r.t. $P$ and is thus not a proper depth measure.
Theorem 7.
$D_{O2}(x)$ takes values in $[0,1]$. Moreover, $D_{O2}(x)=1$ if and only if
$P\{L(x,X_{1},X_{2})\cup L(X_{2},x,X_{1})\cup L(X_{1},X_{2},x)\}=1,$
where $X_{1},X_{2}\sim P$ are i.i.d.
Intuitively, $D_{O2}(x)$ takes a large value if, for random $X_{1},X_{2}$ and the fixed object $x$, one of the three is typically in between the two others. This does not yet guarantee that $x$ is central with respect to $P$ as it is possible that $X_{1}$ (or $X_{2}$) is the in-between point. As such, large values of $D_{O2}(x)$ do not characterize centrality. As an extreme example of this, if $(\mathcal{X},d)$ is the one-dimensional Euclidean space, then $D_{O2}(x)=1$ for all points $x\in\mathbb{R}$. Due to this, we do not pursue the theoretical properties of $D_{O2}$ further. However, we still include $D_{O2}(x)$ in our data examples, where it works surprisingly well in locating the centers of distributions.
4 Simulation examples
4.1 In-sample optimization
Throughout, we abbreviate the depth functions as follows: MOD3 = metric Oja depth, MOD2 = the depth-like counterpart of MOD3 defined in Section 3.3, MHD = metric half-space depth, MLD = metric lens depth, MSD = metric spatial depth. Depths were implemented in Rcpp [11], and are available at https://github.com/vidazamani/Depth-functions-for-Object-Data. All these functions have computational complexity $\mathcal{O}(n^{3})$ for computing the depths of the full sample (assuming [7, Algorithm 1] with the sample itself as anchors is used for MHD), except for MOD3 which is instead $\mathcal{O}(n^{4})$. Note that the complexities of the Euclidean counterparts of these, and other Euclidean depths can be found in [20, Table 1.9].
In our first experiment, we compare the depths with simulations in the context of location estimation. Given a sample $X_{1},\ldots,X_{n}\in\mathcal{X}$ from a distribution $P$, each sample depth function $D_{n}$ is used to determine an in-sample estimate of the deepest object, i.e., $\hat{\mu}:=X_{i_{0}}$ where $i_{0}=\mathrm{argmin}_{i=1,\ldots,n}D_{n}(X_{i})$. The depths are compared based on the averages, over 200 replications, of the estimation errors $d(\hat{\mu},\mu)$ where $\mu\in\mathcal{X}$ is the true location of the distribution $P$. In addition to the error, we evaluate also the computation times of the depths.
4.1.1 Correlation matrix dataset
In our first simulation, we consider the Riemannian manifold $(\mathcal{X},d)$ of $p\times p$ correlation matrices as the data space, where $d(X_{1},X_{2})=\|\mathrm{Log}(X_{1}^{-1/2}X_{2}X_{1}^{-1/2})\|_{F}$ is the affine invariant Riemannian metric, see [2]. Such data are commonly encountered, e.g., in neuroscience in the form of brain connectivity matrices [23].
The random $p\times p$ correlation matrices $X_{1},\ldots,X_{n}$ were generated as $X_{i}:=\mathrm{diag}(S_{i})^{-1/2}S_{i}\mathrm{diag}(S_{i})^{-1/2}$, where $S_{i}:=U_{i}D_{i}U_{i}^{\prime}$, $U_{i}$ is a random orthogonal matrix and $D_{i}$ is a diagonal matrix with all positive diagonal elements, independent of $U_{i}$. The diagonal matrices $D_{i}$ were generated as follows
$D_{i}=\mathrm{diag}\left(\exp(\mathcal{N}(\nu,1)),\underbrace{\exp(\mathcal{N}(-\nu,1)),\ldots,\exp(\mathcal{N}(-\nu,1))}_{p-1\text{ times}}\right).$
The parameter $\nu$ was determined as follows: each $X_{i}$ had the probability $\varepsilon>0$ of being an outlier, in which case $\nu=3$, whereas the remaining $1-\varepsilon$ proportion of the data (the bulk) used $\nu=0$. The matrices $U_{i}$ are uniformly random, see [24] for details. This choice ensures that the eigenvectors of $S_{i}$ do not favor any direction. Consequently, the distribution of $S_{i}$ is rotationally symmetric around the half-line $\{CI_{p}\mid C\geq 0\}$ and, intuitively, the most central matrix w.r.t. this distribution has to reside on this half-line. As the correlation matrices $X_{i}$ are obtained from the $S_{i}$ by scaling the diagonal elements to unity, any reasonable depth measure should estimate the central object of the distribution of $X_{i}$ to be $\mu=I_{p}$. Note that both the bulk and the outliers share the same true most central object (identity matrix) but the high variance of the outliers makes the estimation more difficult for larger $\varepsilon$.
Figure 1: The average estimation errors for each of the five methods in the correlation matrix simulation. The scale of the $y$-axis is logarithmic.
The simulation has three parameters, the contamination proportion $\varepsilon=0.05,0.10$, dimension $p=3,5,10$ and sample size $n=10,20,30,40,50,60$, leading to a total of 36 cases. Note that $n$ was deliberately chosen as low, since object data scenarios typically have much smaller sample sizes than in Euclidean data analysis. The average estimation errors are shown in Figure 1, grouped by $\varepsilon$ and $p$. In all cases, as $n$ increases, the estimation error gradually decreases. Moreover, the complexity of the data space grows as the dimension $p$ of the matrices increases, leading to a decline in accuracy. Increasing the percentage of outliers from $5\%$ to $10\%$ did not lead to significant changes in the results, showing that the performance of the estimators remains stable and indicating that they are indeed robust. In all cases, MOD3 and MOD2 demonstrated the best performance, respectively, while MHD showed the weakest performance. MSD was the third most efficient method and close to MOD3 and MOD2 in performance. From the timing results given in Figure 2, we observe that neither $p$ nor $\varepsilon$ affect the running time. This is because neither has an effect on the size of the sample distance matrix, see also the discussion in Section 6. Additionally, MOD3 and MOD2 have a relatively high time cost, see Figure 2, meaning that these depths essentially offer improved performance at the cost of speed. An alternative version of Figure 1 where the results are shown relative to MSD is presented in Appendix D.
Figure 2: The running times (in seconds) of each method in the correlation matrix simulation. The scale of the $y$-axis is logarithmic.
4.1.2 Hypersphere dataset
In this experiment, we generated samples of points on a $p$-dimensional unit hypersphere $\mathcal{X}$. As a metric $d$ we used the usual arc length distance. Each point $X_{i}$ was simulated by generating the $p$-dimensional random vector $Z_{i}$ from $\mathcal{N}_{p}(\lambda 1_{p},I_{p})$, where $1_{p}$ is a vector of ones and $\lambda\neq 0$, and then taking $X_{i}:=Z_{i}/\|Z_{i}\|$. By symmetry, this strategy leads to the most central object always being equal to $\mu=\mathrm{sign}(\lambda)(1/\sqrt{p})1_{p}$ and the absolute value of the parameter $\lambda$ controls the spread of the points; the larger the value of $|\lambda|$, the more closely the points are concentrated around the location. For the outliers we used $\lambda=-1$ and for the bulk $\lambda=5$, meaning that the two distributions have their most central points on the opposite sides of the hypersphere, see Figure D.1 in Appendix D for an illustration.
Figure 3: The average estimation errors for each of the five methods in the hypersphere simulation. The scale of the $y$-axis is logarithmic.
The simulation thus has three parameters, $\varepsilon$, $p$, $n$, and we use for them the same values as in the earlier simulation. As shown in Figure 3 and its relative version presented in Appendix D, the performance of each metric depth function on the hypersphere dataset is very similar to their performance on the correlation data. We have omitted the timing results for the sphere simulation as they were visually almost identical to Figure 2.
We also performed another simulation in the same setting with the parameters $\varepsilon=0.10$, $n=100,200$, $p=50,100,200,400,800,1600$, where the purpose is to investigate the effect of increasing dimension on the estimate. The average estimation errors over 100 replicates are plotted in Figure 4. From the results we observe that: (i) MSD, MLD, MOD2, MOD3 yield almost identical results and, as expected, improve when $n$ is increased. (ii) MHD is worse than the other four and changing $n$ does seems not affect it. The standard half-space depth is known to degenerate in higher dimensions [10] and, thus, it could be that MHD suffers from the same. A non-trivial solution could be to regularize its computation in a suitable way. (iii) For all methods, the error grows in $p$. From the viewpoint of efficiency, e.g., $n=100,p=500$ yields roughly the same accuracy as $n=200,p=650$. (iv) To obtain a baseline, we considered also an estimator which picks a random object from the sample and uses it as the estimate of the location. This leads, regardless of $p$, to estimation errors of $0.413$, $0.4078$ for $n=100,200$, respectively, showing that all estimators improve considerably upon this naive benchmark.
Figure 4: The effect of increasing the dimension of the unit sphere to the in-sample estimation error, estimated using 100 replications. The lines of MSD, MLD, MOD2 and MOD3 are almost perfectly overlapping.
4.2 Out-of-sample optimization
The in-sample optimization studied in the previous section might come across as a naive approach, but in many metric spaces it is the best one can do, the lack of Euclidean structure preventing the use of standard optimization algorithms. However, despite being non-Euclidean, some metric spaces admit coordinate representations/encodings in Euclidean spaces, e.g., unit spheres (stereographic projection), the positive definite manifold (Cholesky decomposition) and the space of $L_{2}$-functions (Karhunen-Loève expansion, assuming we truncate it). In these cases, the full Euclidean non-linear optimization machinery can be brought forth to find the deepest point.
To formalize the earlier, assume that there exists a bijective map $f:\mathcal{X}\to\mathcal{S}$ where $S\subseteq\mathbb{R}^{q}$ for some $q$. Then, the problem of finding the deepest object of a sample $X_{1},\ldots,X_{n}\in\mathcal{X}$ can be formulated as
$\max_{v\in\mathcal{S}}D_{n}(f^{-1}(v)),$
where $D_{n}$ is some metric depth function w.r.t. the sample and $v$ is an Euclidean vector. After finding the optimizer $v_{0}$, it can be mapped to $\mathcal{X}$ simply as $f(v_{0})$. Since the map $v\mapsto D_{n}(f^{-1}(v))$ is highly non-linear, we propose using bounded Nelder-Mead (NMKB) and Limited-memory BFGS (L-BFGS-B) for the purpose. NMKB is a robust and derivative-free optimizer, which makes it generally well-suited for non-smooth functions like metric depths. However, it can be slow in high-dimensional problems. On the other hand, L-BFGS-B uses approximated gradients to find the optimum solutions and stores only the most recent steps during the optimization process, making it faster and more memory-efficient. See [4] for more details.
We employed here the implementation provided by the R-package dfoptim for NMKB and R’s base optim() function for L-BFGS-B, configuring the algorithm with the following tuning parameters. The initial values were set using the top five in-sample deepest points and the search space was defined with a width of 0.10 for each dimension. To further alleviate computational burden, we apply a dimension reduction via PCA to the representations in the space $\mathcal{S}$ to map them to an $r$-dimensional space with $r\ll q$. As training data for the dimension reduction, we use the images $f(X_{1}),\ldots,f(X_{n})$ of the sample. As data, we generate a sample of correlation matrices similarly as described in Section 4.1.1, meaning that $(\mathcal{X},d)$ is now the manifold of positive definite matrices having unit diagonal. We define the map $f$ such that $f(X_{i})$ is the vector of length $p(p+1)/2$ containing the non-zero elements in the Cholesky decomposition of $X_{i}$. Since the used optimization algorithms can produce solutions lying outside of the set $\mathcal{S}$ (that is, solutions which are covariance matrices, but not correlation matrices), we manually scale the final optimizer $f(v_{0})$ into a proper correlation matrix.
In this simulation, the matrix dimension $p$ was set to 5, and the contamination proportion $\varepsilon$ was fixed to 10%. The sample size, as in previous simulations in Sections 4.1.1 and 4.1.2, was considered to be $n=10,20,30,40,50,60$. A new parameter introduced in this simulation was the number of principal components, determined by a threshold $\texttt{TSH}\in(0,1]$, representing the minimum proportion of total variance to be explained by the chosen principal components. Thus, the number of principal components is $r=\max(2,\{k:\sum_{i=1}^{k}\text{ExplainedVariance}_{i}\geq\text{TSH}\})$. In this simulation TSH was set to be $0.9$. The whole simulation was repeated 500 times using both MHD and MOD3. As further competitors we took the MHD-based in-sample and out-of-sample estimators in the R-package MHD [6], see also [7]. The latter is based on Riemannian optimization of the depth function and denoted by DLP hereafter. Finally, also the in-sample version of MOD3 was included.
Figure 5: Comparing the performance of different optimization methods on MOD3 and MHD. DLP refers to the out-of sample method Left: estimation error of 7 different cases. L-FBSG-B yielded exactly the same result on MOD3 as NMKB (light red line). Right: Estimation Time. The scale of the y-axis is logarithmic.
As observed in the results in Figure 5 (left), the use of the two introduced optimization algorithms (NMKB and L-BFGS-B) combined with dimension reduction has significantly contributed to reducing the estimation error compared to using only the in-sample estimator, which are indicated by the solid lines. What is conspicuous is that, NMKB has the best performance over the two other optimization strategies, DLP and L-BFGS-B. Moreover, MOD3 exhibits lower estimation error than MHD in both out-of-sample and in-sample cases. However, as shown in the Figure 5 (right), MOD3 involves a more time-consuming process compared to MHD. Additionally, despite the substantial computational burden imposed by using NMKB algorithm, we observe that with increasing sample size, even without using NMKB, the in-sample estimators tends to follow a similar trend.
5 Real data example
In this section, we aim to evaluate the performance of the proposed depth functions using real data. The secondary aim of this experiment is to demonstrate how the depth functions, whose analytically complex form prevents theoretical inferential results, can still be used for statistical inference via relying on computational techniques. For this purpose, the Age_pyramids_2014 data from R-package HistDAWass was chosen. This dataset contains histogram-valued data representing the age distribution for each country in the world, for both sexes combined and separately for males and females. We model the histograms as objects in the Wasserstein space $(\mathcal{X},d)$ where $d$ is the $L_{2}$-Wasserstein distance, implemented in the HistDAWass-package.
The current analysis was conducted on a subset of data containing both sexes and all African and European countries, accounting for a total of 95 countries. Based on the geographic locations of the two continents, we hypothesize that age distributions differ between them. This is visually confirmed in Figure 6, where the deepest age distributions, w.r.t. MOD3, among African countries (left) shows a different structure from the corresponding country in Europe (right). Thus, our research question is: Is there a significant difference in the age distribution between these two groups? To study this, we next conduct a permutation test for determining whether the difference in the deepest histograms in the two continents is statistically significant. Letting $f_{1},f_{2}\in\mathcal{X}$ denote the deepest in-sample histograms of the two groups, we use the $L_{2}$-Wasserstein distance between them, $t=d(f_{1},f_{2})$, as the test statistic. To simulate the distribution of $t$ under the null hypothesis (i.e., assuming no difference between the two groups), we randomly permute the group labels of the $n=95$ histograms and compute the test statistic value. This reshuffling was done a total of 500 times, yielding $t^{*}_{1},\ldots,t^{*}_{500}$, using which the $p$-value of the test is computed as $\{t\leq t^{*}_{i}\}/500$.
The resulting $p$-values for the five different depths are extremely small and are shown in Figure 7 (left). All depth functions agree that there is a difference between the groups at the significance level $0.05$.
Figure 6: The deepest age histograms among African (left) and European (right) countries, with respect to $L_{2}$-Wasserstein distance and MOD3. The dashed green lines depict the means of the two distributions.
Figure 7: The $p$-value results of permutation test across depth functions. The test is at level $\alpha$ = 0.05, indicated by the horizontal blue dashed-line.
We next take the previous result of no differences as a ground truth and continue the experiment by (a) contaminating the original data set, and (b) performing the same test for the contaminated data set and observing which of the depth-metric combinations still let us obtain the correct conclusion (no group difference). This experiment therefore mimics the practical scenario where the data has been partially wrongly recorded. To perform the contamination, we randomly select $k$ countries and swap their labels (a European country is labeled an African country and vice versa), using two different choices, $k=12,16$. To reduce randomness in the choice of the swap, we repeat the simulation 10 times, using 100 iterations in the test each time and the resulting average $p$-values are shown in Figure 7 (middle and right). Under medium contamination (12 out of 95 countries mislabeled), all depths continue to yield the correct decision. Whereas, under high contamination (16 out of 95 countries mislabeled) both MHD and MOD2 identify the groups as indistinct. The reason for this is likely that the algorithm for MHD is approximative [7] and MOD2 is not a “true” depth function and does not necessarily measure centrality of its input object, see the discussion after formula (3). The most consistent performance across swaps is given by MOD3 and MSD, which likely relates to the fact that they both use numerical distance information in their computation, unlike MHD and MLD which only use indicators, see Section 6 for more discussion of this. The results also align with the simulation studies in Section 4.1 where MSD was always almost as good as MOD3.
6 Conclusions
As shown in the simulations, MOD3 involves a significant time cost. By using partial U-statistics to randomly choose which elements to include in the triple sum in the sample version of (2), the speed of MOD3 computation could be enhanced. Another important point that could impact future results is the process of selecting and setting appropriate values for the hyper-parameters (take initial values as an example) used in the optimization algorithms.
Our timings in Section 4 did not take into account the computation of the interobject distances since this same step is required by all competing methods. In practice, the full runtime thus comprises of computing the distance matrix and the depths of the sample points. The specifics of the distance computation depend greatly on the metric space in question. For example, on the unit sphere in $\mathbb{R}^{p}$ with the arc length metric, forming the distance matrix is a $O(n^{2}p)$-operation, whereas on the manifold of $p\times p$ positive definite matrices with the Riemannian metric, this operation is of the order $O(n^{2}p^{3})$. The complexity of the depth computation depends only on $n$ and not on other aspects (such as dimensionality) of the sample, in a manner described in the beginning of Section 4.
The computation of MLD and MHD does not rely on the actual distances $d(x_{i},x_{j})$ but only on indicators of the form $\mathbb{I}\{d(x_{i},x_{j})\leq d(x_{i},x_{k})\}$, see Section 2. This means that MLD and MHD (a) are invariant to transformations that preserve these indicators, and (b) can be computed even if one knows only the binary values of these triple comparisons. The latter viewpoint for MLD was taken in [16] to find deepest objects based on binary dissimilarity data. Our MOD3, on the other hand, uses the actual values $d(x_{i},x_{j})$ in its computation, meaning that it cannot be computed based on indicators alone. This limits its usefulness in such dissimilarity scenarios but, at the same time, allows MOD3 to use additional information on the data, which shows as an improved performance in several of our experiments over the other methods. We also note that this behavior is perfectly in line with the original Oja depth, which also depends on the numerical values of the distances $\|x_{i}-x_{j}\|$.
Possible future research topics include: (a) Investigating why $D_{O2}$, despite not being a measure of depth, manages to emprically estimate central objects on the manifold of positive definite matrices and on the unit sphere. (b) Depth functions are typically expected to be maximized at a symmetric center of a distribution. However, unlike, e.g., the metric half-space depth, $D_{O3}$ does not have a natural companion concept of symmetry using which such a center could be defined. One possibility would be to use the quantities $A(x,X_{1},X_{2}):=\{d^{2}(x,X_{1})+d^{2}(x,X_{2})-d^{2}(X_{1},X_{2})\}/\{2d(x,X_{1})d(x,X_{2})\}$, which satisfy $A(x,X_{1},X_{2})\geq-1$ with equality if and only if $x$ lies in between of $X_{1}$ and $X_{2}$ [28, Theorem 1]. This suggests that if there exists an object $x_{0}$ such that
$A(x_{0},X_{1},X_{2})<_{S}A(x,X_{1},X_{2}),\quad\mbox{for all }x\in\mathcal{X}\setminus\{x_{0}\},$ (4)
where $_{S}$ denotes a suitable form of strict stochastic dominance, then $x_{0}$ is a natural candidate for the most central point. Informally, (4) means that $x_{0}$ is more likely to satisfy $d(X_{1},X_{2})\approx d(X_{1},x_{0})+d(x_{0},X_{2})$ than any other point in $\mathcal{X}$. Accordingly, we conjecture that if $x_{0}$ satisfies (4), then it also uniquely maximizes $D_{O3}$.
Acknowledgments
The work of VZ was supported by the Finnish Doctoral Program Network in Artificial Intelligence, AI-DOC (decision number VN/3137/2024-OKM-6). The work of VZ and JV was supported by the Research Council of Finland (Grants 347501, 353769).
Supplementary materials
The supplementary appendices contain proofs of the technical results, additional simulation plots, and other material. The methods and simulation code (in R) are accessible at https://github.com/vidazamani/Depth-functions-for-Object-Data.